A train starts from a station at point A and accelerates uniformly to a speed of 20m/s. It maintains this speed and then retards uniformly until it comes to rest at point B. The distance between the stations is 2000m the total time taken is 2 minutes.if the acceleration is twice the retardation. a) draw a velocity time graph. b) find the time for which the train is moving at constant speed. c) find it's retardation


1
Expert's answer
2022-03-13T18:46:44-0400

Solution.

"v_0=0 m\/s;"

"v=20 m\/s;"

"a=2a_1;"

"t=120s;"

"a)"

"b)\dfrac{v^2}{2a}+\dfrac{v^2}{2a_1}+v\tau=2000;"

"\dfrac{v}{2a}+\dfrac{v}{a_1}+\tau=120;"

"3v^2+4a_1v\tau=8000a_1;"

"3v+2a_1\tau=240a_1;"

"1200+80a_1\tau=8000a_1;"

"60+2a_1\tau=240a_1;" |"\sdot40"

"1200+80a_1\tau=8000a_1;"

"2400+80a_1\tau=9600a_1;"

"-1200=-1600a_1;"

"a_1=0.75 m\/s^2;"

"2a_1\tau=240a_1-3v;"

"\tau=\dfrac{240\sdot0.75-60}{2\sdot0.75}=80s;"

"c)a_1=0.75 m\/s^2;"

Answer: "b)\tau=80s;"

"c)0.75 m\/s^2."